OpenStudy (anonymous):

clear each equation of fractions and solve for z? x x 10 - + - = - 2 6 3

OpenStudy (anonymous):

clear each equation of fractions and solve for x? not z

OpenStudy (anonymous):

\[\frac x2 + \frac x6 = \frac{10}{3}\]\[\frac x2 * \frac33 + \frac x6 = \frac {10}{3} * \frac22\]\[\frac{3x}{6} + \frac x6 = \frac{20}{6}\]\[\frac{4x}{6} = \frac {20}{6}\]\[x = \frac{20}{6} * \frac64\]\[x = \frac{20}{4}\]\[x = 5\]